Journal of Lie Theory Volume 8 (1998) 187 C ◯1998 Heldermann Verlag

Erratum to

“The Quantum Double of a (locally) Compact Group”

T. H. Koornwinder and N. M. Muller

Communicated by Karl H. Hofmann

Abstract. A small error occuring in [1] is corrected.

The classification of conjuagacy classes of SL(2, ℝ) as given in Section 5.2 of [1] is not complete. For trace 2 (cases v. & vi.) there is a third conjugacy class

C_{−1} with representative n_{−1} = ( \begin{matrix}1 & −1 \\ 0 & 1 \end{matrix} ). Similarly, for trace −2 (cases vii. &

viii.), there is a third conjugacy class with representative −n_{−1} . This does not harm the conclusion that SL(2, ℝ) is countably separated, since the continuous

function ( \begin{matrix}a & b \\ c & d \end{matrix} ) ↦ b − c is strictly positive on C_{1} and strictly negative on

C_{−1} .

References

[1] Koornwinder, T. H., and N. M. Muller, The quantum double of a (locally) compact group, Journal of Lie Theory 7 (1997), 101–120.

Department of Mathematics University of Amsterdam Plantage Muidergracht 24 1018 TV Amsterdam The Netherlands thk@wins.uva.nl Institute for Theoretical Physics University of Amsterdam Valckenierstraat 65 1018 XE Amsterdam The Netherlands nmuller@phys.uva.nl

Received August 1, 1997

ISSN 0949–5932 / $2.50 ^{_{C} ◯} Heldermann Verlag